# Groups which are metrically weakly sofic with respect to word norms

### v1

@@ 5.3 Further connections with metric RF
Proposition 5.5 and Corollary 5.2 suggest that Corollary 4.4 can be deduced by proving that $(\mathsf{F}_{2},\parallel\cdot\parallel)$ is not RF. In this section we discuss this way of arguments in a slightly more general situation.
@@
When $\mathsf{F_{n}}^{w}$ is a free group of a group variety with a free base $x_{0},\ldots,x_{i},\ldots$, $i<\mathsf{n}$, where $\mathsf{n}\in\mathbb{N}\cup\{\omega\}$, we preserve the notation above, i.e. we will assume that the set
@@
Sufficiency is obvious. Let us prove necessity. Let $D$ be a finite subset of $(\mathsf{F_{n}}^{w},\parallel\cdot\parallel)$ and $Q$ be a finite subset of $\mathbb{N}$ with $0\in Q$. Let $\varphi:(\mathsf{F_{n}}^{w},\parallel\cdot\parallel)\rightarrow(H,\ell)\in\mathcal{C}$ be as in Definition 2.7. We may assume that there is $D_{0}\subseteq D$ such that $S\cap D=D_{0}$, $D\subseteq\langle D_{0}\rangle$ and for each sub-term $t(\bar{z})$ of $w(\bar{z})$ all values of the form $t(\bar{d})$ with $\bar{d}$ from $D_{0}$ belong to $D$. By quasi-linearity of $w(\bar{z})$ the set $\varphi(D)$ generates a subgroup of $H$ which belongs to $V_{w}$. In particular we can simultaniously extend $\varphi$ and $S\setminus D_{0}\to 1$ to a homomorphism into $H$, say $\hat{\varphi}$. Taking a subgroup if necessary we assume that $\hat{\varphi}$ is surjective.
@@
$\hat{\varphi}(B^{\mathsf{F_{n}}^{w}}_{m}(1))=\hat{\varphi}(\overline{S}^{m})=\hat{\varphi}(\overline{S})^{m}\subseteq B^{H}_{m}(1).$
Indeed, the latter inclusion follows from the triangle inequality. We see that $\ell(\hat{\varphi}(g))\leq\parallel g\parallel$ for any $g\in\mathsf{F_{n}}^{w}$. ∎
We now repeat the argument of Corollary 4.4.

### v2

@@ 5.3 Further connections with metric RF
Proposition 5.5 and Corollary 5.2 suggest that Corollary 4.4 can be deduced by proving that $(\mathsf{F}_{2},\parallel\cdot\parallel)$ is not RF. In this section we discuss this way of argumentation in a slightly more general situation.
@@
When $\mathsf{F}^{w}_{\mathsf{n}}$ is a free group of a group variety with a free base $x_{0},\ldots,x_{i},\ldots$, $i<\mathsf{n}$, where $\mathsf{n}\in\mathbb{N}\cup\{\omega\}$, we preserve the notation above, i.e. we will assume that the set
@@
Sufficiency is obvious. Let us prove necessity. Let $D$ be a finite subset of $(\mathsf{F}^{w}_{\mathsf{n}},\parallel\cdot\parallel)$ and $Q$ be a finite subset of $\mathbb{N}$ with $0\in Q$. Let $\varphi:(\mathsf{F}^{w}_{\mathsf{n}},\parallel\cdot\parallel)\rightarrow(H,\ell)\in\mathcal{C}$ be as in Definition 2.7. We may assume that there is $D_{0}\subseteq D$ such that $S\cap D=D_{0}$, $D\subseteq\langle D_{0}\rangle$ and for each sub-term $t(\bar{z})$ of $w(\bar{z})$ all values of the form $t(\bar{d})$ with $\bar{d}$ from $D_{0}$ belong to $D$. By quasi-linearity of $w(\bar{z})$ the set $\varphi(D)$ generates a subgroup of $H$ which belongs to $V_{w}$. In particular, we can simultaneously extend $\varphi$ and $S\setminus D_{0}\to 1$ to a homomorphism into $H$, say $\hat{\varphi}$. Taking a subgroup if necessary we assume that $\hat{\varphi}$ is surjective.
@@
$\hat{\varphi}(B^{\mathsf{F}^{w}_{\mathsf{n}}}_{m}(1))=\hat{\varphi}(\overline{S}^{m})=\hat{\varphi}(\overline{S})^{m}\subseteq B^{H}_{m}(1).$
Indeed, the latter inclusion follows from the triangle inequality. We see that $\ell(\hat{\varphi}(g))\leq\parallel g\parallel$ for any $g\in\mathsf{F}^{w}_{\mathsf{n}}$. The rest is clear. ∎
We now repeat the argument of Corollary 4.4. This gives a slightly weaker statement, but in a different way.